# Infinite-component $BF$ topological field theory: Nexus of fracton order, Toeplitz braiding, and non-Hermitian Amplification

### v2

@@ III.3 Numerical results
In Fig. 3, we demonstrate the Toeplitz braiding encoded in iBFBF theories with $K$ matrices sharing the same mathematical form of more general Hatano-Nelson Hamiltonians (Eq. (12)), verifying the results obtained from analytic deduction.
Panels (a1)–(e1) demonstrate the braiding statistics encoded in HN-type iBFBF theories with the $K_{\mathrm{HN}}^{\prime}$ matrix specified by $n=5$, $b=1$, $c=5$. (a1) and (b1) are the plots of the singular modes of the corresponding $K_{\mathrm{HN}}^{\prime}$ matrix, and ZSMs are highlighted in red. (c1) is the matrix plot of the braiding phase $\Theta_{I,J}=2\pi({K_{\mathrm{HN}}^{\prime}}^{-1})_{I,J}$ introduced in Eq. (9). The presence of nontrivial braiding phases in the upper-right region of ${K_{\mathrm{HN}}^{\prime}}^{-1}$ indicates that there exist particle excitations located at $w_{1}$ boundary and loop excitations located at $w_{2}$ boundary, which can feel each other via braiding process. In contrast, the absence of the upper-right components of ${K_{\mathrm{HN}}^{\prime}}^{-1}$ shows that the braiding phase vanishes when particle’s and loop’s $w$ coordinates are exchanged. As the # of layers $N$ goes to infinity, the braiding phase $\Theta_{1,N}$ in (d1) oscillates between $(-\pi,\pi]$, while $\Theta_{N,1}$ in (e1) rapidly decays to zero.
Figure 4: (a1) and (a2) display skin modes and eigenvalues of $K_{\mathrm{HN}}^{\prime}$ matrix specified by $n=5,b=5,c=1$. (b1) and (b2) display skin modes and eigenvalues of $K_{\mathrm{HN}}^{\prime}$ matrix specified by $n=5,b=3,c=1$. For illustrative purpose, we depict the skin modes and eigenvalues of the $K_{\mathrm{HN}}^{\prime}$ matrices with system size $N=40$.
@@
Likewise, Panels (a2)–(e2) demonstrate the braiding statistics encoded in NH-type iBFBF theories with the $K_{\mathrm{HN}}^{\prime}$ matrix specified by $n=5$, $b=5$, $c=1$, in which particle excitations located at $w_{2}$ boundary and loop excitations located at $w_{1}$ boundary that can feel each other via braiding process. (a2) and (b2) are the plots of the singular modes of the $K_{\mathrm{HN}}^{\prime}$ matrix, and ZSMs are highlighted in red. (c2) is the matrix plot of the braiding phase $\Theta_{I,J}=2\pi({K_{\mathrm{HN}}^{\prime}}^{-1})_{I,J}$ introduced in Eq. (9). The presence of nontrivial braiding phases in the lower-left region of ${K_{\mathrm{HN}}^{\prime}}^{-1}$ indicates that there exist particle excitations located at $w_{2}$ boundary and loop excitations located at $w_{1}$ boundary, which can feel each other via braiding process. In contrast, the absence of the upper-right components of ${K_{\mathrm{HN}}^{\prime}}^{-1}$ shows that the braiding phase vanishes when particle’s and loop’s $w$ coordinates are exchanged. As the # of layers $N$ goes to infinity, the braiding phase $\Theta_{N,1}$ in (e2) oscillates between $(-\pi,\pi]$, while $\Theta_{1,N}$ in (d2) rapidly decays to zero.
@@
For comparison, we also present an iBFBF theory without Toeplitz braiding in Panels (a3)–(e3). The corresponding $K_{\mathrm{HN}}^{\prime}$ matrix is obtained by the parameters $n=5$, $b=3$, and $c=1$. Panels (a3) and (b3) display the singular modes of the $K_{\mathrm{HN}}^{\prime}$ matrix, where ZSMs are absent. The exponentially suppressed upper-right and lower-left elements in Panel (c3) indicate the absence of nonlocal braiding statistics along the stacking direction, which is further reflected in the asymptotic convergence of $\Theta_{1,N}$ and $\Theta_{N,1}$, as shown in Panels (d3) and (e3).
@@ IV.1 Toeplitz Braiding encoded in Su-Schrieffer-Heeger type $K$ matrix
Case II arises when $|m_{1}|<|n_{1}|$ and $|m_{2}|>|n_{2}|$, leading to $K_{\mathrm{nSSH}}^{-1}$ with exponentially suppressed lower-left elements, while the upper-right elements increase exponentially with the system size. Here, particles at the $w_{1}$ boundary can braid nontrivially with loops at the $w_{2}$ boundary if $n_{2}\,\mathrm{rad}(n_{2})\nmid m_{2}$, while exchanging the $w$ coordinates of the particles and the loops yields trivial statistics. If $n_{2}\,\mathrm{rad}(n_{2})|m_{2}$, the braiding statistics between particles and loops at distinct $w$ boundaries is trivial.
@@ IV.2 ZSMs and Toeplitz braiding
In Case II where $\left\{\begin{array}[]{l}|m_{1}|<|n_{1}|\\ |m_{2}|>|n_{2}|\end{array}\right$., $K_{\mathrm{nSSH}}$ possesses the following LZSM and RZSM, denoted by $\mathbf{u}_{2}$ and $\mathbf{v}_{2}$, respectively.
@@
In line with Eq. (24), we construct an approximate matrix $M_{\mathrm{nSSH}}$ for the inverse $K_{\mathrm{nSSH}}^{-1}$ that captures the contribution of ZSMs to the upper-right or lower-left elements of $K_{\mathrm{nSSH}}^{-1}$. For Case I with $\left\{\begin{array}[]{l}|m_{1}|>|n_{1}|\\ |m_{2}|<|n_{2}|\end{array}\right$., $M_{\mathrm{nSSH}}$ is constructed by $\mathbf{u}_{1}$, $\mathbf{v}_{1}$ and $\sigma_{1}$.
@@
For example, if the system size $N=4$,
@@
For example, if system size $N=4$,
@@
For example, if system size $N=4$,
@@ IV.3 Numerical results
Figure 5: Braiding statistics in different nSSH-type iBFBF theories. (a1)–(e1), (a2)–(e2) and (a3)–(e3) demonstrate the braiding statistics encoded in nSSH-type iBFBF theories with the $K_{\mathrm{nSSH}}$ matrix specified by $\left\{\begin{array}[]{l}n_{1}=n_{2}=2\\ m_{1}=3,m_{2}=1\end{array}\right$., $\left\{\begin{array}[]{l}n_{1}=n_{2}=2\\ m_{1}=1,m_{2}=3\end{array}\right$. and $\left\{\begin{array}[]{l}n_{1}=n_{2}=2\\ m_{1}=3,m_{2}=5\end{array}\right$., respectively. (a1)–(a3) are the plots of the left singular modes of these matrices, while (b1)–(b3) are the plots of the right singular modes of these matrices. The ZSMs are highlighted in red and purple. (c1)–(c3) are the matrix plots of the braiding phase $\Theta_{I,J}=2\pi(K_{\mathrm{nSSH}}^{-1})_{I,J}$, where we take the system size $N=40$ (the resulting matrix size of $K_{\mathrm{nSSH}}$ is $2N=80$) as illustrative examples. (d1)–(d3) and (e1)–(e3) demonstrate how the braiding phases $\Theta_{1,2N}$ and $\Theta_{2N,1}$ vary as the system size $N$ increases.
@@
In Fig. 5, we depict several examples that demonstrate the nonlocal braiding statistics induced by ZSMs, verifying the analytic deductions presented above. Panels (a1)–(e1) demonstrates the braiding statistics encoded in nSSH-type iBFBF theories with the $K_{\mathrm{nSSH}}$ matrix specified by $n_{1}=n_{2}=2$, $m_{1}=3$, $m_{2}=1$, which belongs to Case I. (a1) and (b1) are the plots of the singular modes of the $K_{\mathrm{nSSH}}$ matrix, where ZSMs are highlighted in red. (c1) is the matrix plot of the braiding phase $\Theta_{I,J}=2\pi(K_{\mathrm{nSSH}}^{-1})_{I,J}$ (Eq. (9)). The presence of nontrivial braiding phases in the lower-left region of $\Theta_{I,J}$ indicates that there exist particle excitations located at $w_{2}$ boundary and loop excitations located at $w_{1}$ boundary that can feel each other via braiding process. In contrast, the absence of a nontrivial braiding phase in the upper-right region of $\Theta_{I,J}$ shows that when their $w$ coordinates are exchanged, the braiding phase vanishes. (d1) shows that as the size $2N$ of the matrix increases, the braiding phase $\Theta_{1,2N}$ drastically decays to zero as the system size increases, while (e1) shows that the braiding phase $\Theta_{2N,1}$ oscillates between $(-\pi,\pi]$, reflecting the non-liquid nature of the iBFBF theory.
@@
Likewise, Panels (a2)–(e2) demonstrates the braiding statistics encoded in nSSH-type iBFBF theories with the $K_{\mathrm{nSSH}}$ matrix specified by $n_{1}=n_{2}=2$, $m_{1}=1$, $m_{2}=3$, which belongs to Case II. (a2) and (b2) are the plots of the singular modes of such $K_{\mathrm{nSSH}}$ matrix, where ZSMs are highlighted in red. (c2) is the matrix plot of the braiding phase $\Theta_{I,J}=2\pi(K_{\mathrm{nSSH}}^{-1})_{I,J}$. The presence of nontrivial braiding phases in the upper-right region of $\Theta_{I,J}$ indicates that there exist particle excitations located at $w_{1}$ boundary and loop excitations located at $w_{2}$ boundary that can feel each other via braiding process. In contrast, the absence of a nontrivial braiding phase in the lower-left region of $\Theta_{I,J}$ shows that when their $w$ coordinates are exchanged, the braiding phase vanishes. (d2) shows that as the system size $N$ increases, the braiding phase $\Theta_{1,2N}$ oscillates between $(-\pi,\pi]$, while (e2) shows that the braiding phase $\Theta_{2N,1}$ drastically decays to zero as the system size $N$ increases.
@@
Panels (a3)–(e3) demonstrates the braiding statistics encoded in nSSH-type iBFBF theories with the $K_{\mathrm{nSSH}}$ matrix specified by $n_{1}=n_{2}=2$, $m_{1}=3$, $m_{2}=5$, which belongs to Case III. In this case, the $K_{\mathrm{nSSH}}$ matrix possesses two sets of ZSMs, highlighted in red and purple in Panels (a3) and (b3). Nontrivial braiding statistics arises for both configurations: particles at $w_{1}$ boundary with loops at $w_{2}$ boundary, and vice versa, as is demonstrated in the matrix plot (c3) of the braiding phase $\Theta_{I,J}=2\pi(K_{\mathrm{nSSH}}^{-1})_{I,J}$. (d3) and (e3) show that as the system size $N$ increases, both $\Theta_{1,2N}$ and $\Theta_{2N,1}$ oscillate between $(-\pi,\pi]$.
@@
So far, we focus on the role of ZSMs in nonlocal braiding statistics along the stacking direction. Ref. [110] shows that boundary zero modes obtained from eigendecompostion capture the upper-right and lower-left elements of $K^{-1}$ in the thermodynamic limit when the $K$ matrix is symmetric. Are these two approaches compatible with each other? Are there any differences between the approximate matrix $M$ constructed from ZSMs and the approximate matrix $M^{\prime}$ constructed from boundary zero modes when the $K$ matrix is symmetric? In fact, Eqs. (107) and (114) in Appendix A are equations for the boundary zero modes of the Hermitian $K$ matrices, thus these two approaches are indeed identical. We explicitly verify this using $K_{\mathrm{nSSH}}$ for the case $n_{1}=n_{2}=n$ and $m_{1}=m_{2}=m$. When $n_{1}=n_{2}=n$ and $m_{1}=m_{2}=m$, $K_{\mathrm{nSSH}}$ (Eq. (25)) is a symmetric matrix sharing the same mathematical form of the SSH Hamiltonian [187]. If $|m|>|n|$, the approximate matrix $M_{\mathrm{nSSH}}^{\prime}$ for the inverse $K_{\mathrm{nSSH}}^{-1}$ constructed from boundary zero modes $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ reads [110]

### v3

@@ III.3 Numerical results
In Fig. 3, we demonstrate the Toeplitz braiding encoded in iBFBF theories with $K$ matrices sharing the same mathematical form of more general Hatano-Nelson Hamiltonians [Eq. (12)], verifying the results obtained from analytic deduction.
@@
Figures 3(a1)–3(e1) demonstrate the braiding statistics encoded in HN-type iBFBF theories with the $K_{\mathrm{HN}}^{\prime}$ matrix specified by $n=5$, $b=1$, $c=5$. Figures 3(a1) and 3(b1) are the plots of the singular modes of the corresponding $K_{\mathrm{HN}}^{\prime}$ matrix, and ZSMs are highlighted in red. Figure 3(c1) is the matrix plot of the braiding phase $\Theta_{I,J}=2\pi({K_{\mathrm{HN}}^{\prime}}^{-1})_{I,J}$ introduced in Eq. (9). The presence of nontrivial braiding phases in the upper-right region of ${K_{\mathrm{HN}}^{\prime}}^{-1}$ indicates that there exist particle excitations located at $w_{1}$ boundary and loop excitations located at $w_{2}$ boundary, which can feel each other via braiding process. In contrast, the absence of the lower-left components of ${K_{\mathrm{HN}}^{\prime}}^{-1}$ shows that the braiding phase vanishes when particle’s and loop’s $w$ coordinates are exchanged. As the number of layers $N$ goes to infinity, the braiding phase $\Theta_{1,N}$ in Fig. 3(d1) oscillates between $(-\pi,\pi]$, while $\Theta_{N,1}$ in Fig. 3(e1) rapidly decays to zero.
Figure 4: Panels (a1) and (a2) display skin modes and eigenvalues of $K_{\mathrm{HN}}^{\prime}$ matrix specified by $n=5,b=5,c=1$. Panels (b1) and (b2) display skin modes and eigenvalues of $K_{\mathrm{HN}}^{\prime}$ matrix specified by $n=5,b=3,c=1$. For illustrative purpose, we depict the skin modes and eigenvalues of the $K_{\mathrm{HN}}^{\prime}$ matrices with system size $N=40$.
Likewise, Figs. 3(a2)–3(e2) demonstrate the braiding statistics encoded in NH-type iBFBF theories with the $K_{\mathrm{HN}}^{\prime}$ matrix specified by $n=5$, $b=5$, $c=1$, in which particle excitations located at $w_{2}$ boundary and loop excitations located at $w_{1}$ boundary that can feel each other via braiding process. Figures 3(a2) and 3(b2) are the plots of the singular modes of the $K_{\mathrm{HN}}^{\prime}$ matrix, and ZSMs are highlighted in red. Figure 3(c2) is the matrix plot of the braiding phase $\Theta_{I,J}=2\pi({K_{\mathrm{HN}}^{\prime}}^{-1})_{I,J}$ introduced in Eq. (9). The presence of nontrivial braiding phases in the lower-left region of ${K_{\mathrm{HN}}^{\prime}}^{-1}$ indicates that there exist particle excitations located at $w_{2}$ boundary and loop excitations located at $w_{1}$ boundary, which can feel each other via braiding process. In contrast, the absence of the upper-right components of ${K_{\mathrm{HN}}^{\prime}}^{-1}$ shows that the braiding phase vanishes when particle’s and loop’s $w$ coordinates are exchanged. As the number of layers $N$ goes to infinity, the braiding phase $\Theta_{N,1}$ in Fig. 3(e2) oscillates between $(-\pi,\pi]$, while $\Theta_{1,N}$ in Fig. 3(d2) rapidly decays to zero.
@@
For comparison, we also present an iBFBF theory without Toeplitz braiding in Figs. 3(a3)–3(e3). The corresponding $K_{\mathrm{HN}}^{\prime}$ matrix is obtained by the parameters $n=5$, $b=3$, and $c=1$. Figures 3(a3) and 3(b3) display the singular modes of the $K_{\mathrm{HN}}^{\prime}$ matrix, where ZSMs are absent. The exponentially suppressed upper-right and lower-left elements in Fig. 3(c3) indicate the absence of nonlocal braiding statistics along the stacking direction, which is further reflected in the asymptotic convergence of $\Theta_{1,N}$ and $\Theta_{N,1}$, as shown in Figs. 3(d3) and 3(e3).
@@ IV.1 Toeplitz Braiding encoded in Su-Schrieffer-Heeger type $K$ matrix
Case II arises when $|m_{1}|<|n_{1}|$ and $|m_{2}|>|n_{2}|$, leading to $K_{\mathrm{nSSH}}^{-1}$ with exponentially suppressed lower-left elements, while the upper-right elements increase exponentially with the system size. Here, particles at the $w_{1}$ boundary can braid nontrivially with loops at the $w_{2}$ boundary if $n_{2}\,\mathrm{rad}(n_{2})\nmid m_{2}$, while exchanging the $w$ coordinates of the particles and the loops yields trivial statistics. If $n_{2}\,\mathrm{rad}(n_{2})|m_{2}$, then the braiding statistics between particles and loops at distinct $w$ boundaries is trivial.
@@ IV.2 ZSMs and Toeplitz braiding
In Case II where $\left\{\begin{array}[]{l}|m_{1}|<|n_{1}|\\ |m_{2}|>|n_{2}|\end{array}\right$., $K_{\mathrm{nSSH}}$ possesses the following LZSM and RZSM, denoted by $\mathbf{u}_{2}$ and $\mathbf{v}_{2}$, respectively,
@@
In line with Eq. (24), we construct an approximate matrix $M_{\mathrm{nSSH}}$ for the inverse $K_{\mathrm{nSSH}}^{-1}$ that captures the contribution of ZSMs to the upper-right or lower-left elements of $K_{\mathrm{nSSH}}^{-1}$. For Case I with $\left\{\begin{array}[]{l}|m_{1}|>|n_{1}|\\ |m_{2}|<|n_{2}|\end{array}\right$., $M_{\mathrm{nSSH}}$ is constructed by $\mathbf{u}_{1}$, $\mathbf{v}_{1}$ and $\sigma_{1}$:
@@
For example, if the system size $N=4$, then
@@
For example, if system size $N=4$, then
@@
For example, if system size $N=4$, then
@@ IV.3 Numerical results
Figure 5: Braiding statistics in different nSSH-type iBFBF theories. Panels (a1)–(e1), (a2)–(e2) and (a3)–(e3) demonstrate the braiding statistics encoded in nSSH-type iBFBF theories with the $K_{\mathrm{nSSH}}$ matrix specified by $\left\{\begin{array}[]{l}n_{1}=n_{2}=2\\ m_{1}=3,m_{2}=1\end{array}\right$., $\left\{\begin{array}[]{l}n_{1}=n_{2}=2\\ m_{1}=1,m_{2}=3\end{array}\right$. and $\left\{\begin{array}[]{l}n_{1}=n_{2}=2\\ m_{1}=3,m_{2}=5\end{array}\right$., respectively. Panels (a1)–(a3) are the plots of the left singular modes of these matrices, while panels (b1)–(b3) are the plots of the right singular modes of these matrices. The ZSMs are highlighted in red and purple. Panels (c1)–(c3) are the matrix plots of the braiding phase $\Theta_{I,J}=2\pi(K_{\mathrm{nSSH}}^{-1})_{I,J}$, where we take the system size $N=40$ (the resulting matrix size of $K_{\mathrm{nSSH}}$ is $2N=80$) as illustrative examples. Panels (d1)–(d3) and (e1)–(e3) demonstrate how the braiding phases $\Theta_{1,2N}$ and $\Theta_{2N,1}$ vary as the system size $N$ increases.
@@
In Fig. 5, we depict several examples that demonstrate the nonlocal braiding statistics induced by ZSMs, verifying the analytic deductions presented above. Figures 5(a1)–5(e1) demonstrates the braiding statistics encoded in nSSH-type iBFBF theories with the $K_{\mathrm{nSSH}}$ matrix specified by $n_{1}=n_{2}=2$, $m_{1}=3$, $m_{2}=1$, which belongs to Case I. Figures 5(a1) and 5(b1) are the plots of the singular modes of the $K_{\mathrm{nSSH}}$ matrix, where ZSMs are highlighted in red. Figure 5(c1) is the matrix plot of the braiding phase $\Theta_{I,J}=2\pi(K_{\mathrm{nSSH}}^{-1})_{I,J}$ [Eq. (9)]. The presence of nontrivial braiding phases in the lower-left region of $\Theta_{I,J}$ indicates that there exist particle excitations located at $w_{2}$ boundary and loop excitations located at $w_{1}$ boundary that can feel each other via braiding process. In contrast, the absence of a nontrivial braiding phase in the upper-right region of $\Theta_{I,J}$ shows that when their $w$ coordinates are exchanged, the braiding phase vanishes. Figure 5(d1) shows that as the size $2N$ of the matrix increases, the braiding phase $\Theta_{1,2N}$ drastically decays to zero as the system size increases, while Fig. 5(e1) shows that the braiding phase $\Theta_{2N,1}$ oscillates between $(-\pi,\pi]$, reflecting the non-liquid nature of the iBFBF theory.
@@
Likewise, Figs. 5(a2)–5(e2) demonstrates the braiding statistics encoded in nSSH-type iBFBF theories with the $K_{\mathrm{nSSH}}$ matrix specified by $n_{1}=n_{2}=2$, $m_{1}=1$, $m_{2}=3$, which belongs to Case II. Figures 5(a2) and 5(b2) are the plots of the singular modes of such $K_{\mathrm{nSSH}}$ matrix, where ZSMs are highlighted in red. Figure 5(c2) is the matrix plot of the braiding phase $\Theta_{I,J}=2\pi(K_{\mathrm{nSSH}}^{-1})_{I,J}$. The presence of nontrivial braiding phases in the upper-right region of $\Theta_{I,J}$ indicates that there exist particle excitations located at $w_{1}$ boundary and loop excitations located at $w_{2}$ boundary that can feel each other via braiding process. In contrast, the absence of a nontrivial braiding phase in the lower-left region of $\Theta_{I,J}$ shows that when their $w$ coordinates are exchanged, the braiding phase vanishes. Figure 5(d2) shows that as the system size $N$ increases, the braiding phase $\Theta_{1,2N}$ oscillates between $(-\pi,\pi]$, while FIg. 5(e2) shows that the braiding phase $\Theta_{2N,1}$ drastically decays to zero as the system size $N$ increases.
@@
Figures 5(a3)–5(e3) demonstrates the braiding statistics encoded in nSSH-type iBFBF theories with the $K_{\mathrm{nSSH}}$ matrix specified by $n_{1}=n_{2}=2$, $m_{1}=3$, $m_{2}=5$, which belongs to Case III. In this case, the $K_{\mathrm{nSSH}}$ matrix possesses two sets of ZSMs, highlighted in red and purple in Figs. 5(a3) and 5(b3). Nontrivial braiding statistics arises for both configurations: particles at $w_{1}$ boundary with loops at $w_{2}$ boundary, and vice versa, as is demonstrated in the matrix plot in Figs. 5(c3) of the braiding phase $\Theta_{I,J}=2\pi(K_{\mathrm{nSSH}}^{-1})_{I,J}$. Figures 5(d3) and 5(e3) show that as the system size $N$ increases, both $\Theta_{1,2N}$ and $\Theta_{2N,1}$ oscillate between $(-\pi,\pi]$.
@@
So far, we focus on the role of ZSMs in nonlocal braiding statistics along the stacking direction. Reference [109] shows that boundary zero modes obtained from eigendecompostion capture the upper-right and lower-left elements of $K^{-1}$ in the thermodynamic limit when the $K$ matrix is symmetric. Are these two approaches compatible with each other? Are there any differences between the approximate matrix $M$ constructed from ZSMs and the approximate matrix $M^{\prime}$ constructed from boundary zero modes when the $K$ matrix is symmetric? In fact, Eqs. (107) and (114) in Appendix A are equations for the boundary zero modes of the Hermitian $K$ matrices, thus these two approaches are indeed identical. We explicitly verify this using $K_{\mathrm{nSSH}}$ for the case $n_{1}=n_{2}=n$ and $m_{1}=m_{2}=m$. When $n_{1}=n_{2}=n$ and $m_{1}=m_{2}=m$, $K_{\mathrm{nSSH}}$ [Eq. (25)] is a symmetric matrix sharing the same mathematical form of the SSH Hamiltonian [186]. If $|m|>|n|$, then the approximate matrix $M_{\mathrm{nSSH}}^{\prime}$ for the inverse $K_{\mathrm{nSSH}}^{-1}$ constructed from boundary zero modes $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$ reads [109]